Generate ordered pairs from the function 3x+1 in the domain 0^x^3.
I haven't done ordered pairs yet. Any help or advice would be great. Thank you.

Answer by stanbon(75887)

Generate ordered pairs from the function
f(x) = 3x+1 in the domain 0<= x <=3.
----
If x = 0, y = 3*0+1 = 1 giving you pair (0,1)
---
If x = 1, y = 3*1+1 = 4 giving you pair (1,4)
---
If x = 2, y = 3*2+1 = 7 giving you pair (2,7)
---
If x = 3, y = 3*3+1 = 10 giving you pair (3,10)
==================================================
Cheers,
Stan H.
=============